Charles must make a loan payment of $2,486 at the end of each quarter. His nominal annual loan rate is 2%, compounded quarterly. The original term of the loan was 15 years. How much did Charles originally borrow?
Answer #1

Answer

the originally borrowed amount is equal to the present value of the cash flow

PV=A*(P|A,i,n)

A=2486

i=effective quarterly =nominal annual /number of compounding =2/4=0.5=r

n=number of quarters =15*4=60

PV=2486*(P|A,0.5%,60)

=2486*51.725560751130985=128589.744

the value is $128589.74

======

also, by formula

PV = A4-6+7) PV = 2486 +1-(1+0.005) 60 0.005 PV = 128589.74403